OpenStudy (anonymous):

Give the values of a, b, and c from the general form of the equation (2x + 1)(x - 2) = 0. a) a = 2, b = -3, c = -2 b) a = 2, b = 5, c = -2 c) a = 3, b = 1, c = -1

OpenStudy (radar):

Multiply the expression as indicated: Use FOIL if you wish, You should get the following:

OpenStudy (radar):

\[2x ^{2}-3x-2=0\]

OpenStudy (radar):

a=2, b=-3, c=-2

OpenStudy (anonymous):

you need to use the FOIL method (F) stands for First, (o) stands for Outer, (I) stands for inner and (L) stand for last. So therefore (2xa1)(x-2) so F=2x.x O=2x.-2 I=1.x L=1.-2 So FOIL yeilds 2x^2-4x+X-2 now simplify so we get 2x^2-3x-2. This is a quadratic funtion that has the form ax+bx+c So a=2, b=-3 and c=-2 chow!